OpenStudy (anonymous):

what is the equation of the hyperbola with vertices at (-4,0) and (4,0) and foci at (-6,0) and (6,0)? A. x^2/16 -y^2/20 = 1 B. y^2/16 -x^2/20 = 1 C. x^2/16 -y^2/36 = 1 D. x^2/16 -y^2/52 = 1

jimthompson5910 (jim_thompson5910):

How far did you get here?

OpenStudy (anonymous):

i think i got it is it C?

jimthompson5910 (jim_thompson5910):

That's incorrect

OpenStudy (anonymous):

oh then now i am confused on how to find it out

jimthompson5910 (jim_thompson5910):

What's the center of this hyperbola?

OpenStudy (anonymous):

(0,0)

jimthompson5910 (jim_thompson5910):

How far is it from either vertex?

OpenStudy (anonymous):

4?

jimthompson5910 (jim_thompson5910):

So that's the value of 'a' in the equation below x^2/a^2 - y^2/b^2 = 1 Which means we get this so far x^2/16 - y^2/b^2 = 1

jimthompson5910 (jim_thompson5910):

The focal distance (c) is the length from the center to either focus. This is c = 6 So c^2 = 36 Now use the pythagorean theorem to find b^2 a^2 + b^2 = c^2 16 + b^2 = 36 b^2 = 36 - 16 b^2 = 20

jimthompson5910 (jim_thompson5910):

So x^2/16 - y^2/b^2 = 1 turns into x^2/16 - y^2/20 = 1

jimthompson5910 (jim_thompson5910):

The equation is in the form x^2/a^2 - y^2/b^2 = 1 because it opens right/left (which means that the x^2 term is positive)

OpenStudy (anonymous):

ok thank you that makes sense now...i have two more questions if you can help

jimthompson5910 (jim_thompson5910):

sure, go ahead

OpenStudy (anonymous):

ok i am confused on how to go about this... what are the coordinates of the solution that lies in quadrant 4 of 2x^2+y^2=33 and x^2+y^2+2y=19

jimthompson5910 (jim_thompson5910):

First isolate x^2 in the first equation \[\Large 2x^2+y^2=33\] \[\Large 2x^2=33-y^2\] \[\Large x^2=\frac{33-y^2}{2}\]

jimthompson5910 (jim_thompson5910):

Then plug this into the second equation \[\Large x^2+y^2+2y=19\] \[\Large \frac{33-y^2}{2}+y^2+2y=19\]

jimthompson5910 (jim_thompson5910):

Multiply both sides by 2 to get \[\Large 2\left(\frac{33-y^2}{2}+y^2+2y\right)=2*19\] \[\Large 2\left(\frac{33-y^2}{2}\right)+2y^2+4y=38\] \[\Large 33-y^2+2y^2+4y=38\] \[\Large 33-y^2+2y^2+4y-38=0\] \[\Large y^2+4y-5=0\] I'll let you continue

OpenStudy (anonymous):

ok y=0 or y=1

OpenStudy (anonymous):

i am really confused i am coming up with decimals and i have no idea what i am doing kinda makes me feel dumb cuz im usually always good at math

jimthompson5910 (jim_thompson5910):

y = 0 is incorrect but y = 1 is correct

jimthompson5910 (jim_thompson5910):

either use the quadratic formula or factor

jimthompson5910 (jim_thompson5910):

to factor, think of two numbers that multiply to -5 and add to 4

OpenStudy (anonymous):

-4 and -1

jimthompson5910 (jim_thompson5910):

-4 and -1 multiply to -5?

OpenStudy (anonymous):

sorry i am having a bad math day 5 and -1

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

so that means y^2 + 4y - 5 = 0 turns into (y+5)(y-1)=0

OpenStudy (anonymous):

so plug those in the equations right?

jimthompson5910 (jim_thompson5910):

what do you get for y?

OpenStudy (anonymous):

1 and -5

jimthompson5910 (jim_thompson5910):

we're in the 4th quadrant, so y has to be negative

jimthompson5910 (jim_thompson5910):

that means we only have to care about y = -5

jimthompson5910 (jim_thompson5910):

plug that into any equation involving x and y, then solve for x

OpenStudy (anonymous):

x=2

jimthompson5910 (jim_thompson5910):

Good, the solution as an ordered pair is (2,-5)

OpenStudy (anonymous):

ok now i have another like it i will type it but i wanna solve it and then tell u what i think it is and you tell me if it is correct...what are the coordinates of the solution in the 2nd quadrant of x^+4y^2=100 y=1/32*x^2

jimthompson5910 (jim_thompson5910):

is that a typo in the first equation? is it supposed to be x^2+4y^2=100 ??

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok tell me what you get

OpenStudy (anonymous):

well i got y=2.40 and y=-10.40

jimthompson5910 (jim_thompson5910):

We're in quadrant II, so which y value will we use?

OpenStudy (anonymous):

-10.40

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

no 2.40 sorry

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

use that y value to find x

OpenStudy (anonymous):

x= the sqrt. of -332.8

OpenStudy (anonymous):

wait hold on i think that is wrong

jimthompson5910 (jim_thompson5910):

it should be a real number

OpenStudy (anonymous):